OpenStudy (anonymous):

Please help will medal! Simplify the rational expressions. state and excluded values

OpenStudy (anonymous):

\[\frac{ x+7 }{ x^2+4x-21 }\]

OpenStudy (anonymous):

@cwrw238

OpenStudy (n00bstyle):

Hello alice, have you figured it out already?

OpenStudy (anonymous):

Hello N00bstyle no i havent

OpenStudy (n00bstyle):

Are you familiar with the 'sum-product-rule'?

OpenStudy (anonymous):

no...

OpenStudy (n00bstyle):

If you look at the denominator, you can write that expression differently

OpenStudy (n00bstyle):

Ask yourself the question: what two numbers do I need to add to get '+4', while when I multiply those two numbers I get '-21'

OpenStudy (n00bstyle):

It's like a little puzzle, which you have to solve many times more ;)

OpenStudy (anonymous):

lol thats a nice way to think about it

OpenStudy (n00bstyle):

Let me help you out, try the numbers '-3' and '+7'

OpenStudy (n00bstyle):

When you add those numbers you get +4, and when you multiply them you get -21

OpenStudy (anonymous):

ohhhh i see now

OpenStudy (n00bstyle):

The point is to reformulate the expression in the denominator. So x^2 + 4x - 21 can then be written as: (x - 3) * (x + 7)

OpenStudy (n00bstyle):

Now the formulate can be written as: (x+7)/((x-3)(x+7))

OpenStudy (n00bstyle):

\[\frac{ x+7 }{ (x-3)(x+7) }\]

OpenStudy (n00bstyle):

Now since both the numerator and the denomiator have 'x+7' in it, you can divide that part

OpenStudy (anonymous):

okay do the x+7 cancel out?

OpenStudy (n00bstyle):

this leaves you \[\frac{ 1 }{ x-3 }\]

OpenStudy (n00bstyle):

idd, very good

OpenStudy (anonymous):

c: okay

OpenStudy (anonymous):

thank you

OpenStudy (n00bstyle):

as you probably already know, it is not allowed to divide by zero, so which number will give you an asymptote?

OpenStudy (n00bstyle):

I mean: there is an x in the denominator of the formula we derived

OpenStudy (n00bstyle):

And when that x becomes a value which is equal to a number so the denominator results in being zero, then that value of x will be an asymptote

OpenStudy (n00bstyle):

so in this case the asymptote will be at x = 3

OpenStudy (anonymous):

that makes sence

OpenStudy (n00bstyle):

In the future, see these quadratic functions (x^2 + ... - ...) as little puzzles, and apply the sum-product-rule

OpenStudy (anonymous):

I will

OpenStudy (n00bstyle):

then you have a function written between brackets, and then a lot of times parts will cancel out

OpenStudy (n00bstyle):

good luck

OpenStudy (anonymous):

this was very helpful thank you